# Explainable and externally validated machine learning for neuropsychiatric diagnosis via electrocardiograms

### v1

@@ 4 Discussion
Detecting neuropsychiatric conditions through ECG features may initially seem unconventional, as the ECG is traditionally associated with cardiovascular diagnoses. However, the physiological interplay between the cardiovascular and neuropsychiatric systems offers a unique opportunity for diagnostic innovation. While the mechanisms linking neuropsychiatric conditions to ECG abnormalities remain incompletely understood, some known relations include autonomic nervous system dysfunction, neurodegenerative changes, vascular alterations, electrophysiological disruptions, and hormonal or metabolic factors. For example, Alzheimer’s disease, Parkinson’s disease, and brain damage share common ECG abnormalities, primarily reflecting autonomic dysfunction and brain-heart communication. Across these conditions, reduced heart rate variability (HRV), bradycardia due to parasympathetic dominance, and arrhythmias like atrial fibrillation [20] are frequently observed. Similarly, medication effects, such as those from cholinesterase inhibitors in Alzheimer’s or dopamine agonists in Parkinson’s, can further influence conduction and structural heart changes [21].
@@
Despite these insights, the specific relationship to certain abnormalities remains unclear. Our findings reveal specific ECG patterns as distinctive biomarkers for these conditions, identified using machine learning. Similarly, beyond diagnosis, ECG biomarkers hold promise for therapy management by monitoring treatment responses, risk stratification for medication-related adverse events, early intervention to identify stress or relapse, and longitudinal monitoring for chronic conditions. This interdisciplinary approach bridges neurology, psychiatry, and cardiology, advancing novel diagnostic and therapeutic strategies.
Predictive performance
The exceptional predictive power of specific ECG features highlights their ability to reliably identify neuropsychiatric conditions from a single ECG. Consistently high AUROC scores in both internal and external validations demonstrate the robustness of these features across diverse cohorts. The distinct patterns observed across physiological systems underscore the intricate connection between cardiac and neuropsychiatric health.
Feature importance
@@
In this study, age was identified as a key factor, with older patients contributing more to most conditions. This is consistent with well-documented evidence of brain and structural degeneration associated with aging. For example, previous research has reported abnormal electroencephalograms in patients over 45 years old with conditions such as senile and arteriosclerotic psychosis, involutional psychosis, psychosis with mental deficiency, manic-depressive states, psychoneurosis, and schizophrenia [22].
@@
In terms of neurological conditions, a recent study investigated the association of ECG markers and Alzheimer’s disease and have found high hazard ratios for low QRS duration and for low PR interval [23], which aligns with our findings. Another study reported abormally high P wave axes[24], which also aligns with our findings. For Parkinson’s disease, a study that collected data from various articles and meta-analyses found a clear preponderance of higher incidence and prevalence in males [25], which also aligns with our findings. Similarly, a study conducted on 981 patients demonstrated that high QT interval values are associated with Parkinson’s disease, which again matches our findings. However, the authors report that high QTc interval values were also observed in their cohort [26]. In contrast, our analysis, based on nearly 4,000 samples, suggests that while the raw QT interval is elevated in Parkinson’s patients, the QTc interval, when corrected for heart rate, appears to be lower. This discrepancy may be due to differences in sample size, patient characteristics, or the methodologies used to correct the QT interval, as well as the larger heterogeneity within our cohort.
Regarding psychiatric conditions, left QRS axis deviation have been associated with all-cause dementia patients by a recent study [27], which aligns with low or negative values found in our findings. A similar study correlates with another of our findings, as the authors found that shorter QRS duration is associated with non-Alzheimer dementia [28]. For vascular dementia, prior work identified a significant increase of QTc values on a population group which also match our findings [29]. Finally, the high values of QTc interval for patients with delirium correlates with previous literature work on patients with delirium tremens whose also developed tachyarrhythmias and returned to sinus rhythm after appropriate treatment.
@@
The results presented in this work provide hints on on potentially new ECG markers for neuropsychiatric conditions. This includes high values of RR interval for Alzheimer’s disease, low values of QRS axis for Parkinson’s disease, and low values of QRS duration for vascular dementia. We hypothesize that the high RR interval in Alzheimer’s may reflect altered autonomic regulation due to parasympathetic dysfunction [30]. The low QRS axis in Parkinson’s disease could be linked to basal ganglia dysfunction affecting autonomic control of the heart [31]. Similarly, the low QRS duration in vascular dementia might result from cerebral ischemia and microvascular changes affecting cardiac conduction [32]. These hypotheses warrant further investigation.
Therapy vs therapy-naive confounding
@@
These therapy-related cardiac effects further complicate the distinction between disease progression and treatment influence, underscoring the need for careful monitoring, especially in patients with known cardiac conditions [33]. Future research should aim to stratify newly diagnosed cases from follow-ups, as such differentiation would enable a more precise evaluation of the model’s ability to distinguish between new diagnoses and therapy-induced patterns.
Potential innovations and applications
ECG innovations for neuropsychiatric conditions hold great promise across diagnosis, therapy management, and personalized care. Distinctive ECG patterns can act as biomarkers for more efficient, non-invasive detection of neuropsychiatric disorders. Beyond diagnosis, ECG monitoring contributes to therapy management by tracking treatment responses and identifying early physiological changes related to medication efficacy or adverse effects. Risk stratification also emerges as a vital application, enabling clinicians to predict adverse events such as arrhythmias linked to neuropsychiatric treatments or stress-related cardiovascular complications.
Longitudinal monitoring through periodic or wearable ECG devices provides valuable insights into disease progression, relapse or episodes prediction, facilitating real-time assessments of mental and physical health. ECG-derived markers further support personalized medicine by helping tailor treatments to individual patients. Additionally, early intervention strategies can use subtle ECG changes to detect stress, anxiety, or relapse indicators, ensuring timely clinical responses.
Limitations and future work
ECG is a valuable tool for identifying electrical abnormalities that may correlate with neuropsychiatric conditions; however, it does not provide direct diagnostic confirmation or insights into specific neurophysiological changes underlying these disorders. Many ECG alterations are non-specific and can be attributed to factors such as stress, medication effects, or comorbid conditions, making it challenging to isolate patterns uniquely associated with neuropsychiatric conditions. Additionally, the causal mechanisms linking ECG abnormalities to neuropsychiatric disorders remain poorly understood, highlighting the need for further investigation into these complex relationships.
Future research should explore how ECG abnormalities differ across diverse demographic groups and distinguish them from normal variations, such as those related to age [36]. Investigating the causal connections between ECG features and neuropsychiatric conditions will be essential for advancing our understanding [37]. Further studies should focus on using raw ECG waveforms and validating findings across external datasets to improve diagnostic accuracy [15, 9]. The demonstrated potential of raw ECG waveforms to outperform traditional ECG features emphasizes the importance of refining these approaches to enhance precision and reliability in detecting neuropsychiatric conditions.
@@
The code for dataset preprocessing and reproducing experiments is available in a dedicated code repository.
Author contributions
JMLA, WH, and NS conceived and designed the project. JMLA conducted the full experimental analyses, with WH, and NS supervising them, and WH, EO, and DT providing critical revision of clinical intellectual content. JMLA produced the first draft, the rest of the authors revised it. All authors critically revised the content and approved the final version for publication.
Declaration of interests
The authors declare no competing interests.

### v2

@@ 4.1 Main findings
Detecting neurocognitive disorders through ECG features may initially seem unconventional, as the ECG is traditionally associated with cardiovascular diagnoses. However, the physiological interplay between the cardiovascular and neurocognitive systems offers a unique opportunity for diagnostic innovation. While the mechanisms linking neurocognitive disorders to ECG abnormalities remain incompletely understood, some known relations include autonomic nervous system dysfunction, neurocognitive changes, vascular alterations, electrophysiological disruptions, and hormonal or metabolic factors. For example, Alzheimer’s disease, Parkinson’s disease, and brain damage share common ECG abnormalities, primarily reflecting autonomic dysfunction and brain-heart communication.
@@
Across the investigated disorders, reduced heart rate variability (HRV), bradycardia due to parasympathetic dominance, and arrhythmias like atrial fibrillation [29] are frequently observed. Similarly, medication effects, such as those from cholinesterase inhibitors in Alzheimer’s or dopamine agonists in Parkinson’s, can further influence conduction and structural heart changes [30]. Despite these insights, the specific relationship to certain abnormalities remains unclear. Our findings reveal specific ECG patterns as distinctive features for these disorders, identified using machine learning. Similarly, beyond diagnosis, ECG features hold promise for therapy management by monitoring treatment responses, risk stratification for medication-related adverse events, early intervention to identify stress or relapse, and longitudinal monitoring for chronic disorders. This interdisciplinary approach bridges neurology, psychiatry, and cardiology, advancing novel diagnostic and therapeutic strategies.
@@
The exceptional predictive power of specific ECG features highlights their ability to reliably identify neurocognitive disorders from a single ECG. Consistently high AUROC scores in both internal and external validations demonstrate the robustness of these features across diverse cohorts. The distinct patterns observed across physiological systems underscore the intricate connection between cardiac and neurocognitive health. In this study, age was identified as a key factor, with older patients contributing more to most disorders. This is consistent with well-documented evidence of brain and structural degeneration associated with aging. For example, previous research has reported abnormal electroencephalograms in patients over 45 years old with disorders such as senile and arteriosclerotic psychosis, involutional psychosis, psychosis with mental deficiency, manic-depressive states, psychoneurosis, and schizophrenia [31].
@@
A recent study investigated the association of ECG features and Alzheimer’s disease and have found high hazard ratios for low QRS duration and for low PR interval [32], which aligns with our findings. Another study reported abormally high P wave axes[33], which also aligns with our findings. For Parkinson’s disease, a study that collected data from various articles and meta-analyses found a clear preponderance of higher incidence and prevalence in males [34], which also aligns with our findings. Similarly, a study conducted on 981 patients demonstrated that high QT interval values are associated with Parkinson’s disease, which again matches our findings. However, the authors report that high QTc interval values were also observed in their cohort [35]. In contrast, our analysis, based on nearly 4,000 samples, suggests that while the raw QT interval is elevated in Parkinson’s patients, the QTc interval, when corrected for heart rate, appears to be lower. This discrepancy may be due to differences in sample size, patient characteristics, or the methodologies used to correct the QT interval, as well as the larger heterogeneity within our cohort. Regarding all-cause, left QRS axis deviation have been associated with patients by a recent study [36], which aligns with low or negative values found in our findings. A similar study correlates with another of our findings, as the authors found that shorter QRS duration is associated with non-Alzheimer dementia [15]. For vascular dementia, prior work identified a significant increase of QTc values on a population group which also match our findings [37]. Finally, the high values of QTc interval for patients with delirium correlates with previous literature work on patients with delirium tremens whose also developed tachyarrhythmias and returned to sinus rhythm after appropriate treatment.
From the conditions with high predictive internal performance that can not be externally validated due to missing labels in the external dataset, the high discriminative performance of certain psychiatric and behavioral diagnoses such as eating disorders, substance use disorders, and personality disorders raises important clinical considerations about the physiological manifestations of these conditions. Traditionally, the diagnosis of such disorders relies on behavioral assessments, structured interviews, and, in cases of substance use, confirmatory tests like toxicology screenings. However, the ability of ECG-based models to predict these diagnoses suggests that these conditions may also leave detectable and specific cardi-physiological footprints, possibly through chronic stress, autonomic dysfunction, or cardiotoxic effects of certain substances.
@@
The results presented in this work suggests on potentially new ECG features for neurocognitive disorders. This includes high values of RR interval for Alzheimer’s disease, low values of QRS axis for Parkinson’s disease, and low values of QRS duration for vascular dementia. We hypothesize that the high RR interval in Alzheimer’s may reflect altered autonomic regulation due to parasympathetic dysfunction [38]. The low QRS axis in Parkinson’s disease could be linked to basal ganglia dysfunction affecting autonomic control of the heart [39]. Similarly, the low QRS duration in vascular dementia might result from cerebral ischemia and microvascular changes affecting cardiac conduction [40]. These hypotheses warrant further investigation.
4.2 Limitations
ECG is a valuable tool for identifying electrical abnormalities that may correlate with neurocognitive disorders; however, it does not provide direct diagnostic confirmation or insights into specific neurophysiological changes underlying these disorders. Many ECG alterations are non-specific and can be attributed to factors such as stress, medication effects, or comorbid disorders, making it challenging to isolate patterns uniquely associated with neurocognitive disorders. Additionally, the causal mechanisms linking ECG abnormalities to neurocognitive disorders remain poorly understood, highlighting the need for further investigation into these complex relationships. Nevertheless, we want to raise that previous work also in ECG for diagnoses via ICD-10 in MIMIC-IV has shown that there is none significant label correlations [22], which can be used as a suggestion against confounding effects through co-occurring diseases. Similarly, the work in [15] also suggests the feasibility of finding ECG features abnormalities on incident dementia patients (without treatments) furthing supporting the value of this work.
Future research should explore how ECG abnormalities differ across diverse demographic groups and distinguish them from normal variations, such as those related to age [41]. Investigating the causal connections between ECG features and neurocognitive disorders will be essential for advancing our understanding [42]. Further studies should focus on using raw ECG waveforms and validating findings across external datasets to improve diagnostic accuracy [22, 9]. The demonstrated potential of raw ECG waveforms to outperform traditional ECG features emphasizes the importance of refining these approaches to enhance precision and reliability in detecting neurocognitive disorders.
4.3 Implications
@@
These therapy-related cardiac effects further complicate the distinction between disease progression and treatment influence, underscoring the need for careful monitoring, especially in patients with known cardiac disorders [43]. Future research should aim to stratify newly diagnosed cases from follow-ups, as such differentiation would enable a more precise evaluation of the model’s ability to distinguish between new diagnoses and therapy-induced patterns.
ECG innovations for neurocognitive disorders hold great promise across diagnosis, therapy management, and personalized care. Distinctive ECG patterns can act as features for more efficient, non-invasive detection of neurocognitive disorders. Beyond diagnosis, ECG monitoring contributes to therapy management by tracking treatment responses and identifying early physiological changes related to medication efficacy or adverse effects. Risk stratification also emerges as a vital application, enabling clinicians to predict adverse events such as arrhythmias linked to neurocognitive treatments or stress-related cardiovascular complications.
Longitudinal monitoring through periodic or wearable ECG devices provides valuable insights into disease progression, relapse or episodes prediction, facilitating real-time assessments of mental and physical health. ECG-derived features further support personalized medicine by helping tailor treatments to individual patients. Additionally, early intervention strategies can use subtle ECG changes to detect stress, anxiety, or relapse indicators, ensuring timely clinical responses, see a recent perspective emphasizes that early detection of cognitive impairment in primary care enables earlier interventions, including lifestyle changes and management of comorbid conditions, that can slow cognitive decline and improve overall outcomes for patients and caregivers [46]. Similarly, Recent work highlights the importance of predictive models for neurocognitive disorders, showing that early detection and intervention significantly enhance patient outcomes by enabling timely, tailored therapeutic strategies [47].
5 Statements and declarations
Contributorship
JMLA, WH, and NS conceived and designed the project. JMLA conducted the full experimental analyses, with WH, and NS supervising them, and WH, EO, and DT providing critical revision of clinical intellectual content. JMLA produced the first draft, the rest of the authors revised it. All authors critically revised the content and approved the final version for publication. Computing resources were provided through Carl von Ossietzky University of Oldenburg.
Funding statement
No specific funding was received for this research.
Competing interests
The authors have declared no competing interests.
@@ 5 Statements and declarations
The datasets can be found in raw format in their appropiate source codes https://physionet.org/content/mimic-iv-ecg/1.0/, and http://ecgview.org/. The code for dataset preprocessing and reproducing experiments is available at https://github.com/AI4HealthUOL/CardioDiag.
Ethical statement
This study used publicly available and de-identified datasets. No direct patient interaction or intervention was involved. As these datasets are released under established data use agreements and have been ethically approved for secondary research, additional ethics approval was not sought.
Appendix A Results for additional conditions
In Table LABEL:tab:noext, we show conditions with strong predictive performance on the internal dataset that could not be externally validated due to a lack of corresponding labels on the external dataset. In Table 2, we list conditions with strong internal predictive performance that showed a poor performance on the external dataset. In Figure 1, we show the performance results for the condition G931: Anoxic brain damage.
Table 1: ICD-10 codes of neuropsychiatric disorders with high predictive performance in the internal dataset that are not present in the external dataset and can therefore not be externally validated
ICD-10 Code Description AUROC (Internal)
F5001: norexia nervosa, restricting type. 0.994
F500: Anorexia nervosa. 0.990
F5000: Anorexia nervosa, unspecified. 0.987
F50: Eating disorders. 0.962
F502: Bulimia nervosa. 0.953
F602: Antisocial personality disorder. 0.952
F509: Eating disorder, unspecified. 0.916
F1510: Other stimulant abuse, uncomplicated. 0.903
F1910: Other psychoactive substance abuse, uncomplicated. 0.895
F1220: Cannabis dependence, uncomplicated. 0.891
F151: Other stimulant abuse. 0.886
F1310: Sedative, hypnotic or anxiolytic abuse, uncomplicated. 0.884
F12: Cannabis related disorders. 0.883
F121: Cannabis abuse. 0.879
F91: Conduct disorders. 0.878
F1012: Alcohol abuse with intoxication. 0.873
F149: Cocaine use, unspecified. 0.871
F1210: Cannabis abuse, uncomplicated. 0.870
F13: Sedative, hypnotic, or anxiolytic related disorders. 0.869
F15: Other stimulant related disorders. 0.868
F1023: Alcohol dependence with withdrawal. 0.860
F039: Unspecified dementia. 0.849
F1290: Cannabis use, unspecified, uncomplicated. 0.846
F129: Cannabis use, unspecified. 0.845
F1022: Alcohol dependence with intoxication. 0.845
F0390: Unspecified dementia, unspecified severity… 0.843
F1490: Cocaine use, unspecified, uncomplicated. 0.843
F603: Borderline personality disorder. 0.840
F250: Schizoaffective disorder, bipolar type. 0.834
F068: Other specified mental disorders due to known physiological condition. 0.834
F11: Opioid related disorders. 0.833
F132: Sedative, hypnotic or anxiolytic-related dependence. 0.833
F1123: Opioid dependence with withdrawal. 0.829
F112: Opioid dependence. 0.823
F0391: Unspecified dementia with behavioral disturbance. 0.821
F111: Opioid abuse. 0.820
F0280: Dementia in other diseases classified elsewhere, unspecified severity… 0.820
F1120: Opioid dependence, uncomplicated. 0.819
F141: Cocaine abuse. 0.818
F1110: Opioid abuse, uncomplicated. 0.816
F1993: Other psychoactive substance use, unspecified with withdrawal. 0.815
F102: Alcohol dependence. 0.811
F1920: Other psychoactive substance dependence, uncomplicated. 0.810
F1190: Opioid use, unspecified, uncomplicated. 0.808
F14: Cocaine related disorders. 0.803
F181: Inhalant abuse. 0.801
F1810: Inhalant abuse, uncomplicated. 0.801
F1410: Cocaine abuse, uncomplicated. 0.800
F18: Inhalant related disorders. 0.800
F909: Attention-deficit hyperactivity disorder, unspecified type. 0.797
F1020: Alcohol dependence, uncomplicated. 0.795
F98: Other behavioral and emotional disorders… 0.795
F02: Dementia in other diseases classified elsewhere. 0.789
F028: Dementia in other diseases classified elsewhere. 0.789
F205: Residual schizophrenia. 0.788
F192: Other psychoactive substance dependence. 0.787
F90: Attention-deficit hyperactivity disorders. 0.783
F10: Alcohol related disorders. 0.782
F0150: Vascular dementia, unspecified severity, without disturbance. 0.777
F1320: Sedative, hypnotic or anxiolytic dependence, uncomplicated. 0.775
F0151: Vascular dementia with behavioral disturbance. 0.773
F1420: Cocaine dependence, uncomplicated. 0.767
F4322: Adjustment disorder with anxiety. 0.767
F131: Sedative, hypnotic or anxiolytic-related abuse. 0.754
F609: Personality disorder, unspecified. 0.752
F60: Specific personality disorders. 0.751
F259: Schizoaffective disorder, unspecified. 0.749
F1121: Opioid dependence, in remission. 0.745
F101: Alcohol abuse. 0.743
F119: Opioid use, unspecified. 0.743
F39: Unspecified mood [affective] disorder. 0.734
F25: Schizoaffective disorders. 0.732
F429: Obsessive-compulsive disorder, unspecified. 0.732
F122: Cannabis dependence. 0.729
F1011: Alcohol abuse, in remission. 0.727
F1010: Alcohol abuse, uncomplicated. 0.725
F142: Cocaine dependence. 0.723
F191: Other psychoactive substance abuse. 0.720
F42: Obsessive-compulsive disorder. 0.720
F109: Alcohol use, unspecified. 0.712
F1721: Nicotine dependence, cigarettes. 0.710
F332: Major depressive disorder, recurrent severe without psychotic markers. 0.702
F1021: Alcohol dependence, in remission. 0.701
ICD-10 Code Description AUROC (Internal) AUROC (External)
F311: Bipolar disorder, current episode manic without psychotic markers. 0.715 0.628
F322: Major depressive disorder, single episode, severe without psychotic markers. 0.7956 0.520
G931: Anoxic brain damage 0.717 0.800
Table 2: ICD-10 codes of neuropsychiatric disorders with internal AUROC above 0.7 that does not have at least 0.7 in external or does not fall into the neurocognitive disorders category
Figure 1: Performance in terms of AUROC, calibration curve, net benefit, as well as explainability for G931: Anoxic brain damage.
Appendix B Feature comparison across binary outcomes
ICD-10 Code: Description Feature MIMIC-IV Negative MIMIC-IV Positive ECG-VIEW II Negative ECG-VIEW II Positive
G30: Alzheimer’s disease $R$ 769.00 (645.00, 909.00) 769.00 (638.00, 909.00) 857.00 (741.00, 968.00) 870.00 (750.00, 984.00)
PR $\uparrow$ 158.00 (140.00, 178.00) 162.00 (142.00, 185.00) 158.00 (144.00, 172.00) 162.00 (148.00, 180.00)
QRS 94.00 (85.00, 107.00) 94.00 (84.00, 114.00) 90.00 (84.00, 98.00) 88.00 (80.00, 96.00)
QT $\uparrow$ 394.00 (360.00, 428.00) 400.00 (364.00, 436.00) 392.00 (368.00, 416.00) 404.00 (378.00, 430.00)
QTc $\uparrow$ 447.44 (426.03, 473.33) 455.69 (434.61, 482.02) 421.00 (405.00, 442.00) 435.00 (417.00, 454.00)
P wave axis $\uparrow$ 51.00 (32.00, 64.00) 52.50 (32.00, 67.00) 53.00 (37.00, 65.00) 54.00 (38.00, 66.00)
QRS axis $\downarrow$ 14.00 (-15.00, 46.00) 0.00 (-31.00, 31.00) 48.00 (19.00, 68.00) 36.00 (10.00, 55.00)
T wave axis $\uparrow$ 42.00 (13.00, 71.00) 48.00 (12.00, 81.00) 44.00 (27.00, 60.00) 50.00 (32.00, 64.00)
G20: Parkinson’s disease $R$ 769.00 (645.00, 909.00) 800.00 (667.00, 923.00) 857.00 (741.00, 968.00) 800.00 (690.00, 909.00)
PR $\uparrow$ 158.00 (140.00, 178.00) 164.00 (144.00, 187.00) 158.00 (144.00, 172.00) 160.00 (146.00, 178.00)
QRS 94.00 (85.00, 107.00) 96.00 (86.00, 113.00) 90.00 (84.00, 98.00) 86.00 (80.00, 96.00)
QT 394.00 (360.00, 428.00) 400.00 (366.00, 434.00) 392.00 (368.00, 416.00) 390.00 (364.00, 416.00)
QTc $\uparrow$ 447.65 (426.08, 473.38) 447.82 (426.56, 474.31) 421.00 (405.00, 442.00) 433.00 (411.00, 457.00)
P wave axis $\downarrow$ 51.00 (32.00, 64.00) 48.00 (26.00, 64.00) 53.00 (37.00, 65.00) 52.00 (34.00, 67.00)
QRS axis $\downarrow$ 14.00 (-15.00, 46.00) 0.00 (-30.00, 29.00) 48.00 (19.00, 68.00) 25.00 (-2.00, 51.00)
T wave axis $\uparrow$ 42.00 (13.00, 71.00) 48.00 (15.00, 77.00) 44.00 (27.00, 60.00) 49.00 (29.00, 66.00)
F01: Vascular dementia $R$ 769.00 (645.00, 909.00) 769.00 (652.00, 909.00) 857.00 (741.00, 968.00) 779.00 (667.00, 909.00)
PR $\uparrow$ 158.00 (140.00, 178.00) 164.00 (144.00, 188.00) 158.00 (144.00, 172.00) 164.00 (146.00, 182.00)
QRS 94.00 (85.00, 107.00) 96.00 (86.00, 120.00) 90.00 (84.00, 98.00) 88.00 (82.00, 98.00)
QT $\uparrow$ 394.00 (360.00, 428.00) 404.00 (368.00, 439.00) 392.00 (368.00, 416.00) 396.00 (366.00, 428.00)
QTc $\uparrow$ 447.44 (426.04, 473.33) 460.13 (437.80, 484.80) 421.00 (405.00, 442.00) 446.00 (426.00, 467.00)
P wave axis 51.00 (32.00, 64.00) 51.00 (27.50, 65.00) 53.00 (37.00, 65.00) 51.00 (33.00, 64.00)
QRS axis $\downarrow$ 14.00 (-15.00, 46.00) 0.00 (-33.00, 32.00) 48.00 (19.00, 68.00) 33.00 (5.00, 60.00)
T wave axis $\uparrow$ 42.00 (13.00, 71.00) 54.00 (15.00, 95.00) 44.00 (27.00, 60.00) 53.00 (29.00, 78.00)
F03: Dementia $R$ $\downarrow$ 769.00 (645.00, 909.00) 759.00 (638.00, 909.00) 857.00 (741.00, 968.00) 800.00 (682.00, 938.00)
PR $\uparrow$ 158.00 (140.00, 178.00) 164.00 (142.00, 188.00) 158.00 (144.00, 172.00) 160.00 (144.00, 176.00)
QRS 94.00 (85.00, 107.00) 97.00 (86.00, 121.00) 90.00 (84.00, 98.00) 86.00 (80.00, 94.00)
QT $\uparrow$ 394.00 (360.00, 428.00) 400.00 (364.00, 437.00) 392.00 (368.00, 416.00) 394.00 (366.00, 424.00)
QTc $\uparrow$ 447.21 (426.01, 473.12) 459.20 (436.36, 485.76) 421.00 (405.00, 442.00) 437.00 (415.00, 460.00)
P wave axis 51.00 (32.00, 64.00) 51.00 (27.00, 67.00) 53.00 (37.00, 65.00) 52.00 (35.00, 65.00)
QRS axis $\downarrow$ 14.00 (-14.00, 47.00) -2.00 (-32.00, 30.00) 48.00 (19.00, 68.00) 27.00 (3.00, 54.00)
T wave axis 42.00 (14.00, 71.00) 50.00 (12.00, 87.00) 44.00 (27.00, 60.00) 44.00 (24.75, 67.00)
F05: Delirium (physiological) $R$ $\downarrow$ 769.00 (645.00, 909.00) 714.00 (594.00, 845.00) 857.00 (741.00, 968.00) 723.00 (600.00, 857.00)
PR $\downarrow$ 158.00 (140.00, 178.00) 154.00 (136.00, 178.00) 158.00 (144.00, 172.00) 156.00 (140.00, 174.00)
QRS 94.00 (85.00, 107.00) 96.00 (86.00, 116.00) 90.00 (84.00, 98.00) 88.00 (80.00, 98.00)
QT $\downarrow$ 394.00 (360.00, 428.00) 386.00 (348.00, 424.00) 392.00 (368.00, 416.00) 380.00 (348.00, 416.00)
QTc $\uparrow$ 447.21 (426.00, 473.08) 458.39 (435.04, 485.35) 421.00 (405.00, 442.00) 448.00 (422.00, 472.00)
P wave axis 51.00 (32.00, 64.00) 52.00 (31.00, 66.00) 53.00 (37.00, 65.00) 52.00 (36.00, 67.00)
QRS axis $\downarrow$ 14.00 (-15.00, 46.00) 6.00 (-24.00, 41.00) 48.00 (19.00, 68.00) 35.00 (4.00, 63.00)
T wave axis $\uparrow$ 42.00 (14.00, 70.00) 48.00 (8.00, 85.00) 44.00 (27.00, 60.00) 57.00 (31.00, 87.50)
Table 3: Comparison of ECG-derived feature medians and interquantile ranges (IQR) between patients with and without neurocognitive diagnoses in the MIMIC-IV and ECG-VIEW II cohorts. Neurocognitive disorders analyzed include Alzheimer’s disease (G30), and Parkinson’s disease (G20), vascular dementia (F01), unspecified dementia (F03), and delirium due to physiological disorders (F05). Feature distributions are reported separately for positive (with diagnosis) and negative (without diagnosis) cases in each cohort. To highlight consistent trends across both cohorts, an uparrow $\uparrow$ indicates features with increased values in positive cases, while $\downarrow$ indicates features with decreased values
Table 3 presents median values with interquartile ranges for various ECG features stratified by diagnosis codes related to neurocognitive disorders across two datasets. For each ICD-10 code, values are compared between patients without (negative) and with (positive) diagnosis. Notably, for Alzheimer’s disease, the positive samples against the negative ones show an increase in PR, QTc, P wave axis, T wave axis, and a decrease in QRS axis. For Parkinson’s disease, the positive samples against the negative shows an increase of PR, T wave axis, as well as a decrease of P wave axis, QRS axis. For vascular dementia, positive samples against negative show an increase in the PR, QT, QTc, T wave axis, as well as a decrease in the QRS axis. For unspecified dementia, the positive samples against the negative shows an increase of PR, QT, QTc, as well as a decrease of RR, QRS axis. Finally, for delirium, the positive samples against the negative shows increase of QTc, T wave axis, as well as decrease of RR, PR, QT, QRS axis. Overall, the table highlights subtle but potentially clinically relevant ECG feature alterations associated with the investigated diagnoses, which may contribute to improved cardiac monitoring or risk stratification in these populations.